# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v1

@@ 5.2.4 Benchmark models
Tab. LABEL:t:benchmark_models_environment presents the benchmark models employed in the ”Environment” group by at least two articles examined in this review.
Table 7: List of benchmark models in the ”Environment” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_28], [art_33], [art_34], [art_38]
@@
Naive [art_29], [art_34]
Vector autoregression (VAR) [art_28], [art_34], [art_38]
Classical machine learning methods Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_33], [art_46]
Combined FC-LSTM and convolution neural network (CFCC-LSTM) [CFCC-LSTM] [art_35], [art_39]
@@
Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_4], [art_41]
Convolutional long-short term memory network (ConvLSTM) [ConvLSTM] [art_28], [art_41]
Feed-forward neural network (FNN) [art_2], [art_7], [art_38]
@@
Gated recurrent unit (GRU) [art_2], [art_6], [art_7], [art_8], [art_38], [art_43], [art_46]
Informer [Informer] [art_30], [art_47]
Long-short term memory network (LSTM) [art_1], [art_2], [art_3], [art_4], [art_5], [art_6], [art_7], [art_8], [art_28], [art_30], [art_32], [art_33], [art_35], [art_36], [art_38], [art_39], [art_43], [art_46], [art_47]
@@
Support vector regression (SVR) [art_3], [art_6], [art_35], [art_36], [art_38]
Temporal convolutional network (TCN) [TCN] [art_6], [art_30]
@@
Hybrid wavelet random forest and deep belief network (Wavelet-DBN-RF) [art_31], [art_40], [art_44]
Extreme gradient boosting (XGBoost) [art_1], [art_2], [art_5], [art_46]
GNN methods Deeper graph convolutional network (DeeperGCN) [DeeperGCN] [art_1], [art_5]
Graph convolutional neural network and long-short term memory (GC-LSTM) [GC-LSTM] [art_1], [art_2], [art_5], [art_6], [art_7]
Graph convolutional network (GCN) [art_38], [art_41]
Graph WaveNet (GWN) [GWN] [art_1], [art_6], [art_44], [art_47]
@@
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_6], [art_32], [art_44]
@@
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_3], [art_6], [art_28], [art_35], [art_47]
@@
The majority of the benchmarks used fall within the category of classical machine learning, although there are several GNN benchmarks as well. The two most common benchmark models are LSTM and GRU. These models are widely used by environmental data forecasters due to their ability to capture temporal dependencies, and they are considered a reliable benchmark in the field. As for GNN models, the most common benckmarks are the graph convolutional neural network and long-short term memory (GC-LSTM) [GC-LSTM] (which appears only in quality air forecasting problems) and spatio-temporal graph convolutional network (STGCN) [STGCN]. The GC-LSTM is a model that integrates graph convolutional networks and long-short-term memory networks, proposed by Qi et al. in 2019 to forecast the spatio-temporal variation of PM2.5 concentration. The STGCN model instead was proposed in the traffic domain by Yu et al. in 2018. This model is composed of two spatio-temporal convolutional blocks and a fully-connected output layer. Each convolutional block contains two temporal gated convolution layers and one spatial graph convolution layer in the middle. The code for the STGCN model is available at https://github.com/VeritasYin/STGCN_IJCAI-18. Finally, it is important to point out that the multivariate time series forecasting with graph neural network (MTGNN) model proposed by Wu et al. in 2020 in Ref. [MTGNN] is often the second-best performing model in the papers using it, thus it deserves more attention and further investigation in future research.
@@ 5.2.5 Results
In each paper, the proposed model significantly improves on the benchmarks. The accuracy of the models is measured in terms of MAE, RMSE, and MAPE. Among the papers, there is a high diversity in the choice of the proposed datasets. However, there are two sets of papers using the same dataset, whose results can be compared.
The first dataset is the CCMP wind speed dataset (available at http://data.remss.com), which contains 3,942 geographical grid points from the east and southeast sea areas of China. The three papers using it ([art_31], [art_40], [art_44]) select 120 points from the area and use data records from January 2010 to April 2019 at 6 hours intervals, with a total of 13,624 samples at each node, for 10 different forecasting horizons. All papers present results for the complete set of 120 nodes and a subset of nodes. However, since their results are similar, only the average results for all the 120 nodes are reported in Tab. LABEL:t:risultati_ccmp for the sake of simplicity.
Table 8: Comparison of the average accuracy of the different models on the CCMP wind speed dataset for time horizons from 6 hours to 7 days ahead, expressed in terms of MAE and RMSE. The smallest errors are underlined. Numbers from the papers.
@@
For all the time horizons, the most recent dynamic adaptive spatio-temporal graph neural network (DASTGN) model [art_44] has the highest accuracy. Other two well performing models are the data-based adaptive graph learning network (DAGLN) [art_31] and the multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
The second dataset shared by two papers ([art_1], [art_5]) is the one of China National Urban Air Quality (available at https://github.com/Friger/GAGNN). It contains air quality index data (AQI) of 209 cities, collected from January 2017 to April 2019 at 1 hour granularity, for a total of 20,370 samples. Both the papers also integrate exogenous weather variables (humidity, wind direction, wind speed, rainfall, air pressure, temperature, and visibility) collected from Envicloud (http://www.envicloud.cn/). The average accuracy of the models used in the two papers is compared in Tab. LABEL:t:risultati_chinese_aqi.
Table 9: Comparison of the average accuracy of the different models on the Chinese AQI dataset for time horizons from 1 to 6 hours ahead, expressed in terms of MAE and RMSE. The smallest errors are underlined. Numbers from the papers.
@@
The hybrid interpretable neural network and graph neural network model (INNGNN) [art_5] is the most accurate model for all time horizons except 1 hour, where it is beaten by the meta-graph convolutional recurrent network model (MegaCRN) [MegaCRN]. The authors attribute this result to the fact that in the short term the AQI is strongly influenced by nearby cities, and the MegaCRN model is more focused on local features.
@@ 5.3 Finance
”Finance” is another emerging field of study, though not yet widely explored, with 9 out of 156 selected papers. The difficulty of predicting financial data is a well-established fact. This is due to the inherently complex nature of markets, the influence of geopolitical events on them, and the unpredictable behavior of humans which is often irrational and difficult to predict. This complexity makes financial time series highly volatile, and accurate forecasts and data modeling are essential to develop effective trading strategies. GNNs are taken into consideration to model the inter-dependencies between variables together with series dynamics.
@@ 5.3.1 Overview
The ”Finance” group is relatively narrow, in part due to the limited number of published papers. The majority of studies are concentrated on stock price prediction ([art_59], [art_60], [art_62], [art_63], [art_65]), with the aim of forecasting the future trend of the price of a company. Sometimes the stock prediction problem is seen as a classification task. Indeed, due to the stochastic nature and volatility of financial markets, predicting price movements (”up”, ”neutral”, ”down”) can be more feasible than accurately forecasting specific prices. For example, assuming a daily granularity, a rise (”up”) can be considered to occur when the next trading day’s closing price ($p_{t}$) is at least 0.55% higher than the previous day’s closing price ($p_{t-1}$), whereas if $p_{t}$ is more than 0.50% lower than $p_{t-1}$, the price movement can be classified as a ”down”. Otherwise, it is classified as ”neutral”. All the selected papers addressing stock prices approach the problem from a classification perspective. Another distinctive aspect in this group of papers is that researchers usually also evaluate how their predictions affect the stability of investment returns over a specified period. This is commonly done through trading simulations in the test set, evaluated with metrics such as the Sharpe ratio and overall returns. These simulations require modeling rather than forecasting.
An examination of the selected papers indicates an interest in the application of GNNs in Finance from 2022. Two of the papers were published in IEEE Access by IEEE, while the remaining ones appeared in various other journals. Most of the corresponding authors are affiliated with academic institutions in China.
@@ 5.3.2 Datasets
Table 10: List of public datasets in the ”Finance” group and their corresponding links.
@@
China A-Shares [art_63], [art_65], [art_66] https://finance.sina.com.cn/stock
@@
CSI 300 index [art_59], [art_62] https://global.csmar.com/
@@
All the listed datasets represent collections of stock indices. The most common datasets include the China Securities Index (CSI), the 100 and 500 Standard and Poor’s (S&P) indices, and the 100 and 300 China A-Shares indices. Many studies also include exogenous data, such as other market data, events, and news. As for the news, they can be processed by some language model like BERT (as in Refs. [art_59] and [art_66]). Most of the data have a granularity of 1 day (which is typical of financial data), and the forecasting horizon is also generally 1 day. Only one of the selected papers (Ref. [art_62]) analyses also longer time horizons for stock forecasting. As for pre-processing, half of the selected papers use normalization techniques, including min-max normalization, Z-score normalization, and logarithmic normalization.
5.3.3 Proposed models
In terms of the taxonomy of the proposed GNN models, 5 out of 9 papers utilize attentional GNNs, while the other 4 papers employ convolutional GNNs. As for the graph structure, some papers use models which learn the graph structure themselves ([art_62], [art_64]), while some others defines it a priori based on the relationships between companies and stocks ([art_63], [art_67], [art_68]).
Since the majority of papers deal with a stock movement classification problem, the most common loss function is the cross entropy. The majority of the papers explicitly indicate that Python is the programming language utilized, with nearly half employing PyTorch and a few utilizing TensorFlow. Only Ref. [art_63] provides a link to the source code of the proposed knowledge graph and graph convolution neural network (KG-GCN) model, available at https://github.com/Gjl12321/KG_GCN_Stock_Price_Trend_Prediction_System.

### v2

@@ 5.2.3 Types of models
Dynamic Spatial-Temporal Graph Convolutional Recurrent Network (DSTGCRN) [art_360] https://github.com/PalaceTony/DSTGCRN
Dynamic spatiotemporal graph neural network (DST-GNN) [art_354] https://github.com/liaohaibing/MGST_GNN/tree/master
@@
GNNCoder [art_356] https://doi.org/10.5281/zenodo.13358007
GroundwaterFlowGNN [art_359] https://github.com/mlttac/GroundwaterFlowGNN
@@
Physics-aware graph networks (PaGN) [art_365] https://github.com/eunviho/novel_pagn
@@ 5.2.4 Benchmark models
Tab. LABEL:t:benchmark_models_environment presents the benchmark models employed in the “Environment” group by at least two articles examined in this review.
Table 11: List of benchmark models in the “Environment” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_28], [art_33], [art_34], [art_38], [art_363], [art_370], [art_371]
@@
Naive [art_29], [art_34], [art_366]
Vector autoregression (VAR) [art_28], [art_34], [art_38], [art_360]
Non-GNN machine learning methods Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_33], [art_46]
Combined FC-LSTM and convolution neural network (CFCC-LSTM) [CFCC-LSTM] [art_35], [art_39], [art_370]
DLinear [DLinear] [art_31], [art_364]
@@
Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_4], [art_41], [art_368], [art_371]
Convolutional long-short term memory network (ConvLSTM) [ConvLSTM] [art_28], [art_41], [art_355]
Feed-forward neural network (FNN) [art_2], [art_7], [art_38], [art_360], [art_365], [art_370]
@@
Gated recurrent unit (GRU) [art_2], [art_6], [art_7], [art_8], [art_38], [art_43], [art_46], [art_361], [art_364], [art_365], [art_367], [art_370]
Informer [Informer] [art_30], [art_47], [art_354], [art_364], [art_371]
Long-short term time series network (LSTNet) [LSTNet] [art_355], [art_373]
@@
Random forest (RF) [art_6], [art_364], [art_370]
Support vector regression (SVR) [art_3], [art_6], [art_35], [art_36], [art_38], [art_360], [art_370], [art_371]
Recurrent neural network (RNN) [art_46], [art_365], [art_367]
Temporal convolutional network (TCN) [TCN] [art_6], [art_30], [art_355], [art_356], [art_368], [art_371], [art_373]
@@
Hybrid wavelet random forest and deep belief network (Wavelet-DBN-RF) [Wavelet-DBN-RF] [art_31], [art_40], [art_44]
Extreme gradient boosting (XGBoost) [art_1], [art_2], [art_5], [art_46], [art_364]
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_6], [art_360], [art_365], [art_368]
Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_363], [art_368]
Deeper graph convolutional network (DeeperGCN) [DeeperGCN] [art_1], [art_5]
Graph convolutional neural network and long-short term memory (GC-LSTM) [GC-LSTM] [art_1], [art_2], [art_5], [art_6], [art_7], [art_354], [art_361]
Graph convolutional network (GCN) [art_38], [art_41], [art_357], [art_367]
Graph WaveNet (GWN) [GWN] [art_1], [art_6], [art_44], [art_47], [art_363], [art_368]
@@
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_6], [art_32], [art_44], [art_355], [art_360], [art_373]
@@
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_3], [art_6], [art_28], [art_35], [art_47], [art_363], [art_368], [art_371]
@@
Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_368], [art_371]
@@
The majority of the benchmarks used in this thematic group fall within the category of non-GNN machine learning methods, although there are several GNN benchmarks as well. The two most common benchmark models are LSTMs and GRUs. These models are widely used by environmental data forecasters due to their ability to capture temporal dependencies, and are considered a reliable benchmark in the field. As for GNN models, the most common benckmarks are the graph convolutional neural network and long-short term memory newtork (GC-LSTM) [GC-LSTM] (which appears only in quality air forecasting problems), and the spatio-temporal graph convolutional network (STGCN) [STGCN]. The GC-LSTM network is a model that integrates graph convolutional networks and long-short-term memory networks. It was proposed by Qi et al. in 2019 to forecast the spatio-temporal variation of PM2.5 concentration. In the original paper, the GC-LSTM model works with a pre-defined graph based on the geographical distance between air quality stations. The input layer of the GC-LSTM model receives historical air quality observations, meteorological variables, spatial predictors, and temporal predictors. This information is processed through a graph convolution operation to extract spatial features. After that, the temporal feature extraction is performed using a LSTM network, which takes as input the graph convolutional features concatenated with the original signals. The STGCN model, on its part, was proposed in the traffic domain by Yu et al. in 2018, and works with pre-defined adjacency matrices computed based on the distances among stations in the traffic network. The STGCN model is composed of two spatio-temporal convolutional blocks and a fully-connected output layer. Each convolutional block contains two temporal gated convolution layers and one spatial graph convolution layer in the middle. The code for the STGCN model is available at https://github.com/VeritasYin/STGCN_IJCAI-18. Finally, it is important to point out that the so-called multivariate time series forecasting with graph neural network (MTGNN) model proposed by Wu et al. in 2020 in Ref. [MTGNN] is often reported in the papers using it as the second-best performing model, thus it probably deserves special attention and further investigation in future research.
@@ 5.2.5 Results
In each paper of this group, it is stated that the proposed model is significantly better than the benchmarks. The accuracy of the models is measured in terms of MAE, RMSE, and MAPE. Among the papers, there is a high diversity in the choice of the proposed datasets. However, there are two sets of papers using the same dataset, whose results are compared hereafter.
The first dataset is the CCMP wind speed dataset (available at http://data.remss.com), which contains 3,942 geographical grid points from the east and southeast sea areas of China. The three papers using it (Refs. [art_31, art_40, art_44]) select 120 points from the area and use data records from January 2010 to April 2019 at 6 hours intervals, with a total of 13,624 samples at each node, for 10 different forecasting horizons. All papers present results for the complete set of 120 nodes and a subset of nodes. Since these results are similar, only the average results for all the 120 nodes are reported in Tab. LABEL:t:risultati_ccmp for the sake of simplicity.
Table 12: Comparison of the average accuracy of the different models on the CCMP wind speed dataset for different time horizons, expressed in terms of MAE and RMSE. The smallest errors are underlined. Numbers from the original papers.
@@
For all the time horizons considered in the papers, the dynamic adaptive spatio-temporal graph neural network (DASTGN) model [art_44] proposed by Gao et al. in 2023 has the highest accuracy. The DASTGN model is composed of a graph generation module and a spatio-temporal convolution module. The dynamic adaptive graph generation module generates a static graph to model long-term stable dependencies, and a dynamic graph to model the short-term dynamics. These graphs are then merged into an optimal dynamic graph. The spatio-temporal convolution module consists of an input module, a stack of spatio-temporal layers connected with residual and skip connections, and an output module that generates wind speed forecasts. Other two well performing models are the data-based adaptive graph learning network (DAGLN) [art_31], and the multivariate time series forecasting with graph neural network (MTGNN) [MTGNN]. The DAGLN model was proposed by Ren at al. in 2023, and uses an encoder-decodes architecture. The encoder includes an adaptive graph learning module that learns the optimal adjacency matrix starting from a pre-defined one obtained by using the DTW algorithm, a graph convolution with node adaptive parameter learning algorithm for personalized learning of parameter space, and a GRU for temporal modeling. Then, this information is fed into the decoder that operates a convolutional and a linear operation to produce the wind speed forecast. The MTGNN model, proposed by Wu et al. in 2020, is a well-known state-of-the-art model for multivariate time series forecasting. It incorporates a graph learning layer that learns a graph adjacency matrix from scratch, followed by a stack of graph convolution and temporal convolution modules to capture spatial and temporal dependencies, and an output module. To avoid the problem of vanishing gradients, residual and skip connections are added between spatial and temporal convolutional modules.
The second dataset shared by two papers (Refs. [art_1, art_5]) is called China National Urban Air Quality dataset (available at https://github.com/Friger/GAGNN). It contains air quality index data (AQI) of 209 cities, collected from January 2017 to April 2019 at 1 hour granularity, for a total of 20,370 samples. Both these papers integrate exogenous weather variables (humidity, wind direction, wind speed, rainfall, air pressure, temperature, and visibility) collected from Envicloud (http://www.envicloud.cn/). The average accuracy of the models used in the two papers is compared in Tab. LABEL:t:risultati_chinese_aqi.
Table 13: Comparison of the average accuracy of the different models on the Chinese AQI dataset for time horizons from 1 to 6 hours ahead, expressed in terms of MAE and RMSE. The smallest errors are underlined. Numbers from the original papers.
@@
As from Tab. LABEL:t:risultati_chinese_aqi, the hybrid interpretable neural network and graph neural network model (INNGNN) [art_5] is the most accurate model for all time horizons except the 1 hour horizon, where it is beaten by the meta-graph convolutional recurrent network model (MegaCRN) [MegaCRN]. The authors attribute this result to the fact that in the short term the AQI is strongly influenced by nearby cities, whereas the MegaCRN model is more focused on local features. The INNGNN model, proposed by Ding et al. in 2023, consists of a temporal and a spatial dependency module. The temporal module employs self-attention and interpretable neural networks (INN) to extract local and global features of time series data, effectively identifying periodicity and trend of the AQI time series. After that, the spatial module has an encoder-decoder architecture where GNN modules extract spatial dependencies across geographic locations, represented by an adjacency matrix based on spatial distance. The MegaCRN model, introduced by Jiang et al. in 2023 for traffic forecasting, includes a meta-graph learner module and an encoder-decoder spatial module built using a series of convolutional GNNs. The source code for the MegaCRN model can be accessed at https://github.com/deepkashiwa20/MegaCRN.
@@ 5.3 Finance
“Finance” is another emerging theme of study, though not yet widely explored, with just 16 out of 281 selected papers. The difficulty of predicting financial data is a well-established fact. This is due to the inherently complex nature of markets, the influence of geopolitical events on them, and the unpredictable behavior of humans which is often irrational and difficult to predict. This complexity makes financial time series highly volatile, and accurate forecasts and data modeling are essential to develop effective trading strategies. Spatio-temporal GNNs are taken into consideration to model the inter-dependencies between variables together with series dynamics.
@@ 5.3.1 Overview
The “Finance” thematic group is relatively narrow in terms of sub-themes. This is in part due to the limited number of published papers. The majority of studies are concentrated on stock price forecasting [art_59, art_60, art_62, art_63, art_65, art_377, art_382], with the aim of forecasting the future trend of the price of a company. Sometimes the stock prediction problem is seen as a classification task. This is because, due to the stochastic nature and volatility of financial markets, predicting just the direction of price movements (“up”, “neutral”, “down”) can be sometimes more feasible than accurately forecasting price values. In this case, assuming a daily granularity, a rise (“up”) can be considered to occur when the next trading day’s closing price ($p_{t}$) is at least 0.55% higher than the previous day’s closing price ($p_{t-1}$), whereas if $p_{t}$ is more than 0.50% lower than $p_{t-1}$, the price movement can be classified as a “down”. Otherwise, it is classified as “neutral”. All the selected GNN papers addressing stock prices indeed approach the price forecasting problem from a classification perspective. Another distinctive aspect in this group of papers is that researchers usually evaluate, among others, how their predictions affect the stability of investment returns over a specified period. This is commonly done through trading simulations in the test set, evaluated with metrics such as the Sharpe ratio and overall returns [art_59, art_60, art_65, art_66, art_377]. These simulations require modeling rather than forecasting.
An examination of the selected papers indicates an interest in the application of spatio-temporal GNNs in Finance which starts from 2022. Two of the papers were published in IEEE Access by IEEE, while the remaining ones appeared in various other journals. Most of the corresponding authors are affiliated with academic institutions in China.
@@ 5.3.2 Datasets
Table 14: List of public datasets in the “Finance” group and their corresponding links.
@@
China A-Shares [art_63], [art_65], [art_66], [art_377], [art_378] https://finance.sina.com.cn/stock
Criptocurriences [art_381] https://coinmarketcap.com/
@@
CSI 300 index [art_59], [art_62], [art_382] https://global.csmar.com/
Gold and silver market [art_374] https://www.kitco.com/
HAL agriculture [art_379] www.hal.gov.tr
@@
US crude oil market [art_374] http://www.eia.gov/
All the listed datasets represent collections of stock indices. The most common datasets include the China Securities Index (CSI), the 100 and 500 Standard and Poor’s (S&P) indices, and the 100 and 300 China A-Shares indices. Many studies also include exogenous data, such as other market data, events, and news. News can be processed by some language model like BERT (as in Refs. [art_59, art_66, art_378]). Most of the data have a granularity of 1 day, and the forecasting horizon is also generally 1 day. Two of the selected papers (Refs. [art_62, art_377]) analyses longer time horizons for stock forecasting, and others do the same for agricultural price (Ref. [art_379]) or cryptocurrencies’ price (Ref. [art_381]). As for pre-processing techniques, half of the selected papers use normalization techniques, including min-max normalization, Z-score normalization, and logarithmic normalization.
5.3.3 Types of models
In terms of the taxonomy of the proposed GNN models, 6 out of 16 papers utilize attentional GNNs, 9 papers employ convolutional GNNs, and one paper takes into account both convolutional and attentional GNNs (Ref. [art_374]). As for the graph structure, some papers use models which learn the graph structure themselves (e.g., Refs. [art_62, art_64, art_378, art_379, art_382]), whereas some others define it a priori, based on the relationships between companies and stocks (e.g., Refs. [art_63, art_67, art_68, art_381]).
Half of the papers deal with a stock movement classification problem, so the most commonly used loss function is cross entropy. The majority of the papers explicitly indicate that Python is the programming language utilized, with nearly half employing PyTorch, and a few utilizing TensorFlow. Only Refs. [art_63] and [art_379] provide a link to the source code of the proposed models, which are, respectively, the knowledge graph with graph convolution neural network (KG-GCN) model, available at https://github.com/Gjl12321/KG_GCN_Stock_Price_Trend_Prediction_System, and the spectral temporal graph neural network (StemGNN), available at https://github.com/cevher/Time-Series-Analysis.